OpenStudy (anonymous):

How do I work this problem: If x/4=y/5, then x+4/4=?/5 ? Step by step please.

OpenStudy (anonymous):

This doesnt make much sense

OpenStudy (anonymous):

x = 4y/5 x + 4/4 = 4y/5+4/4= (4y+4*5)/4*5=4(y +5)/4*5=y+5/5